OpenStudy (anonymous):

Someone Help Me with Impulse! I got the Impulse but I don't know how to apply it to problem solving, I mean I do but I am not good if any tough question or a little complicated question is asked that requires Impulse concept. I fail to solve it :(

OpenStudy (anonymous):

By impulse do u mean impulse response in signals and systems?

OpenStudy (anonymous):

Umm No I meant the impulse in Newtonian Mechanics \[\vec{J} = \vec{p_2} - \vec{p_1}\] \[\vec{p_i} = m \times \vec{v_{i}}\]

OpenStudy (anonymous):

Impulse is simply another word for change in momentum. Thus, for one object, we can say\[\vec J = \Delta \vec p =m\Delta \vec v = m\left(\vec v_f - \vec v_i\right)\]Remember also that since \(\vec F = \frac{\Delta \vec p}{\Delta t}\) we can also express the following:\[\vec F \Delta t = \Delta \vec p = \vec J\]If you are familiar with calculus, for time-variant force, this is also the following.\[\vec J = \int \vec F\ dt\] If you haven't worked with integral, din't worry about that last one. Give me the problem that is troubling you and I can give you some pointers.

OpenStudy (anonymous):

hahaha....same problem wid me...

OpenStudy (anonymous):

ultimate cases of impulse in my clsrum package

OpenStudy (anonymous):

this cant be solved if one doesnt knw how to use it